# Newtonian potential from scattering amplitudes in super-renormalizable gravity

### v1

@@ Abstract
Using scattering amplitudes, we compute the coupling between a general super-renormalizable gravity theory and massive scalar particles. This allows us to derive the $D$-dimensional Newtonian potential at both tree-level and one-loop level—the latter containing the first calculation by using newly derived three-graviton Feynman rules. In four-dimensional spacetime, we numerically demonstrate that the Newtonian potential remains finite at the origin, providing compelling evidence that the singularity-free nature of super-renormalizable gravity persists at the one-loop level.
Introduction. The quest to reconcile quantum mechanics with general relativity (GR) is not merely a technical issue of taming ultraviolet (UV) divergences, but must also address the fate of the classical singularities. A powerful bridge to this quantum-classical interface is provided by the effective field theory approach, where quantum field theory (QFT) techniques are deployed to extract classical gravitational dynamics from quantum amplitudes Duff:1973zz ; Bjerrum-Bohr:2018xdl ; Goldberger:2004jt ; Kosower:2018adc ; Cheung:2018wkq ; Bern:2019nnu ; Chung:2019yfs ; Donoghue:1993eb ; Bjerrum-Bohr:2002fji ; Donoghue:1994dn ; Donoghue:1993eb ; Bjerrum-Bohr:2002fji ; Neill:2013wsa ; Damour:2016gwp . In this framework, GR emerges from the exchange of spin-2 gravitons Donoghue:1994dn , and modern on-shell methods, such as scattering amplitudes and generalized unitarity, have dramatically refined our ability to reveal the deeper connections between QFT and spacetime dynamics Bern:2008qj ; Mangano:1990by ; Aoude:2023fdm ; DiVecchia:2021bdo ; Brandhuber:2021eyq ; Monteiro:2021ztt ; Manohar:2022dea ; Aoude:2022trd ; Menezes:2022tcs ; Kalin:2022hph ; Guevara:2019fsj ; Aoude:2021oqj ; Herrmann:2021lqe ; Cristofoli:2020uzm .
Scattering amplitudes provide a uniquely powerful framework for probing quantum gravity (QG), with gauge-invariant observables laying the foundation for a future QG phenomenology Kalin:2020fhe ; Bern:2021yeh ; Bern:2021dqo ; Adamo:2022qci ; Bohnenblust:2024hkw ; Jakobsen:2021lvp ; Mougiakakos:2024nku . Their theoretical precision enables the systematic extraction of QG effects, including the constraining of higher-derivative operators induced by loop divergences in effective approaches Wang:2015jna . While tree-level amplitudes already encode rich physical content, their direct computation is often obscured by gauge redundancies and field-redefinition ambiguities in the off-shell Lagrangian formulation DeWitt:1967uc ; Berends:1974gk . To overcome these challenges, modern amplitude methods-most notably generalized unitarity-have been developed, which isolate the physical, “cut-constructible” parts of loop amplitudes by focusing on their non-analytic structure Bern:1994zx . This approach has proven transformative, enabling efficient calculations of long-range interactions in gravity that are governed precisely by such non-analytic terms, scaling as $m/\sqrt{-q^{2}}$ for a massive probe $m$ and momentum transfer $q$ Donoghue:1993eb . These technical advances now drive concrete progress in classical GR, facilitating high-precision computations of post-Newtonian and post-Minkowskian dynamics in binary systems Blanchet:2006jqj .
@@
Super-renormalizable gravity provides a consistent and predictive framework for ultraviolet-complete QG, preserving unitarity, covariance, and perturbative control across all scales Biswas:2011ar ; Modesto:2011kw ; Briscese:2013lna ; Tomboulis:2015esa ; Calcagni:2014vxa ; Alexander:2012aw ; Dona:2015tra . Its defining feature is weak nonlocality—a property shared by string theory Brekke:1987ptq —which ensures super-renormalizability while excluding pathological states. Although traditional perturbative studies in super-renormalizable gravity have established a finite Newtonian potential at $\mathcal{O}(G)$ order Biswas:2011ar and explored other physical aspects Briscese:2013lna ; Krasnikov:1987yj ; Modesto:2011kw ; Briscese:2013lna ; Tomboulis:2015esa ; Dona:2015tra , the behavior at $\mathcal{O}(G^{2})$ order has remained a critical open question, we resolve this outstanding challenge by applying modern amplitude techniques to perform the first complete one-loop computation of the Newtonian potential.
We focus here on a general super-renormalizable gravity scenario described by the action Krasnikov:1987yj ; Biswas:2011ar ; Modesto:2011kw ; Briscese:2013lna ; Tomboulis:2015esa ; Calcagni:2014vxa ; Alexander:2012aw ; Dona:2015tra
@@
$S=\frac{2}{\kappa^{2}}\int\sqrt{-g}\,\mathrm{d}^{D}x\left(R+RF_{1}(\Box)R+R_{\mu\nu}F_{2}(\Box)R^{\mu\nu}+R_{\mu\nu\alpha\beta}F_{4}(\Box)R^{\mu\nu\alpha\beta}\right),$ (1)
where $R$ denotes the Ricci scalar curvature and $\kappa^{2}=32\pi G_{N}$. The functions of the covariant d’Alembertian operator $F_{i}(\Box)$, called form factors, are assumed to be entire functions and can therefore be expanded in a Taylor series as $F_{i}(\Box)=\sum^{\infty}_{n=0}f_{in}\frac{\Box^{n}}{M^{2n}_{\ast}}$ ($i=1,2$), where $M_{\ast}$ denotes the mass scale at which the higher-derivative terms in the action become relevant. The form factors $F_{i}(\Box)$ can be expressed in terms of exponentials of entire functions $H_{\ell}(\Box)$ (with $\ell=0,2$) Koshelev:2017ebj .
@@
$\begin{split}&F_{1}(\Box)=-\frac{(D-2)\left(e^{H_{0}(\Box)}-1\right)+D\left(e^{H_{2}(\Box)}-1\right)}{4(D-1)\Box}+F_{4}(\Box),\\ &F_{2}(\Box)=\frac{e^{H_{2}(\Box)}-1}{\Box}-4F_{4}(\Box).\end{split}$ (2)
The requirement of super-renormalizability imposes two key constraints. First, the form factor $F_{4}(\Box)$ must exhibit the same UV asymptotic behavior as $F_{i}(\Box)$ ($i=0,2$), a condition most straightforwardly satisfied by setting $F_{4}(\Box)=0$, thereby reducing to two independent form factors. Second, the polynomials defining $H_{0}(\Box)$ and $H_{2}(\Box)$ must share the same degree, while the entire functions $e^{H_{i}(\Box)}$ must simultaneously fulfill three distinct consistency conditions Tomboulis:2015esa .
Newtonian potential at tree-level. The classical limit of QFT, corresponding to $\hbar\to 0$, is conventionally identified with the stationary-point approximation $\delta S=0$. Nevertheless, Feynman diagrams to all loop orders can still contribute classically, as comprehensively discussed in Kosower:2018adc and specifically in the gravitational context Bjerrum-Bohr:2018xdl ; Cheung:2020gyp ; Bjerrum-Bohr:2019kec . In this framework, we analyze the vertex function for massive scalars interacting with gravitons: the scalars are treated as point particles with finite momenta, while gravitons carry soft momenta that vanish in the classical limit. This scaling applies to both external and loop momenta, implying that the classical limit corresponds physically to the long-range limit of the interaction. The vertex rules come from the expansion of the matter term of the action which is

### v2

@@ Abstract
Using scattering amplitudes, we compute the coupling between a general super-renormalizable gravity and massive scalar particles. This allows us to derive the $D$-dimensional Newtonian potential at both tree-level and one-loop level—the latter containing the first calculation by using newly derived three-graviton Feynman rules. In four-dimensional spacetime, we numerically demonstrate that the Newtonian potential remains finite at the origin, providing compelling evidence that the singularity-free nature of super-renormalizable gravity persists at the one-loop level.
I Introduction
The challenge of formulating a consistent theory of quantum gravity (QG) has persisted for decades, with various approaches proposed in an effort to reconcile quantum mechanics with general relativity (GR). One of the central concerns in QG is the issue of renormalizability, particularly in the ultraviolet (UV) regime. This is a critical region, as it is where both quantum and classical singularities manifest. Unfortunately, GR proves problematic in the UV limit, suffering from divergences that render it non-renormalizable Utiyama and DeWitt (1962); Buchbinder et al. (1992); Shapiro (2008). However, effective quantum gravity in the classical limit offers a viable framework for recovering GR, where quantum field theory (QFT) techniques are employed to derive classical results from a quantum perspective Duff (1973); Bjerrum-Bohr et al. (2018); Goldberger and Rothstein (2006); Kosower et al. (2019); Cheung et al. (2018); Bern et al. (2019b, a); Cristofoli et al. (2019); Chung et al. (2020). Within this framework, gravitational interactions are treated as mediated by spin-2 gravitons, and GR is recast in the formalism of QFT Donoghue (1995). In addition, the analytical techniques developed in QFT are crucial for refining theoretical predictions, particularly in the study of gravitational waves. In this context, modern approaches such as on-shell scattering amplitudes and generalized unitarity have proven especially effective. While quantum corrections to gravity remain negligible in current experimental settings, they continue to be of great interest in the ongoing research into QG. These methods offer new insights into gravitational phenomena, potentially revealing deeper connections between quantum mechanics and spacetime dynamics.
Scattering amplitudes provide a uniquely powerful framework for probing QG, with gauge-invariant observables laying the foundation for a future QG phenomenology Kälin et al. (2020); Bern et al. (2022, 2021); Adamo et al. (2023); Bohnenblust et al. (2025); Jakobsen et al. (2022); Mougiakakos and Vanhove (2024). Their theoretical precision enables the systematic extraction of QG effects, including the constraining of higher-derivative operators induced by loop divergences in effective approaches Wang and Yin (2015). It is well-established that even at the tree-level, intermediate computational steps often display a complex structure that typically cancels out in the final physical amplitudes DeWitt (1967); Berends and Gastmans (1975); Elvang and Huang (2013). This complexity arises from the inherent ambiguity of the off-shell Lagrangian formalism, which is invariant under gauge transformations and field redefinitions. Recently, there has been a surge of interest in developing advanced methods for simplifying the calculation of scattering amplitudes. The primary methods employed in this context are unitarity techniques, which isolate the ”cut-constructible” components of loop amplitudes, specifically targeting the non-analytic terms Bern et al. (1994). The technique has significantly streamlined amplitude calculations, making many of the current computations for the Standard Model in particle physics feasible, where they would otherwise have been intractable. In classical gravity, the long-range contributions of interest are precisely of this non-analytic form, depending on the dimensionless ratio $\frac{m}{\sqrt{-q^{2}}}$, where $m$ represents a massive probe and $q$ is the appropriately defined momentum Donoghue (1994). This observation has prompted the suggestion that these modern methods be employed to compute post-Newtonian and post-Minkowskian corrections within GR, particularly in the context of astrophysical phenomena such as binary black hole mergers Blanchet (2006); Fraser and Eardley (2015); Minazzoli and Chauvineau (2009); Abbott and others (2016). The application of these advanced computational tools is expected to enhance the precision of theoretical models and provide valuable insights into gravitational wave physics and other relativistic phenomena.
@@ I Introduction
Super-renormalizable gravity offers a consistent and predictive framework for ultraviolet-complete QG, preserving unitarity, covariance, and perturbative control across all energy scales Biswas et al. (2012); Modesto (2012); Briscese et al. (2014); Tomboulis (2015); Calcagni and Modesto (2015); Alexander et al. (2012); Donà et al. (2015). A defining feature of this approach is weak nonlocality—a property shared by string theory Brekke et al. (1988)—which ensures super-renormalizability while effectively eliminating pathological states. While traditional perturbative studies in super-renormalizable gravity have successfully established a finite Newtonian potential at $\mathcal{O}(G)$ Biswas et al. (2012) and explored various phenomenological implications Briscese et al. (2014); Krasnikov (1987); Modesto (2012); Tomboulis (2015); Donà et al. (2015), the theory’s behavior at $\mathcal{O}(G^{2})$ has remained a critical open question. In this work, we address this outstanding challenge by leveraging modern scattering amplitude techniques to perform the first complete one-loop computation of the Newtonian potential.
The paper is organized as follows: In Section II, we introduce the super-renormalizable gravity from the perspectives of unitarity and power counting. In Section III, we give the relation between the scattering amplitudes and the matter fields in the process of gravitational radiation from a massive scalar particle, and subsequently obtained the general expression for the higher-order metric correction. Using this method, we calculate the metric correction and Newtonian potential in $D$-dimensional spacetime at the tree-level. In Section IV, by constructing the Feynman rules for the three-graviton vertex, we investigate the loop-level correction to the Newtonian potential. We demonstrate that the potential retains well-defined singularity even at the one-loop order. In Section V, where we also draw our conclusions.
II Super-renormalizable Gravity
Due to the intrinsic limitations of GR, a variety of modified gravity models have emerged naturally over time as attempts to extend or replace the GR, each aiming to address unresolved issues such as the nature of dark energy, singularities, and the behavior of gravity at small scales. These models include a range of modifications to the Einstein-Hilbert action, from higher-order derivatives of the curvature tensor to nonlocal corrections that introduce additional degrees of freedom (d.o.f) to the gravitational field Nojiri et al. (2017); Nojiri and Odintsov (2011, 2006); Capozziello and De Laurentis (2011); Faraoni and Capozziello (2011); de la Cruz-Dombriz and Saez-Gomez (2012); Olmo (2011). Among these modifications, string-inspired nonlocal gravity has attracted significant attention due to their promising quantum properties and their potential to provide a more robust framework for understanding gravity at quantum scales Deser and Woodard (2007). Initially, the nonlocal gravity was primarily proposed as a possible explanation for the accelerated expansion of the universe. However, the model quickly evolved to provide a broader framework for addressing a variety of quantum gravity phenomena Bamba et al. (2012); Zhang and Sasaki (2012); Elizalde et al. (2012); Nojiri et al. (2011). The profound impact of nonlocal gravity is particularly evident in their connection to string theory, where nonlocal terms are commonly introduced to describe string interactions at high energies. Additionally, the model offers a fresh perspective on the mathematical structure of gravity, with modifications to the gravitational action that go beyond traditional local interaction, introducing both theoretical elegance and physical insights into the UV regime Nojiri and Odintsov (2008); Zhang et al. (2016).
To describe physical phenomena at both classical and quantum levels, most nonlocal models introduce modifications involving either nonlocal scalar fields or terms that involve the d’Alembertian operator $\Box$. These terms allow for a nonlocal dependence of the gravitational field on spacetime, in contrast to the local interactions present in GR. Without losing generality, we focus here on a general nonlocal model described by the following action
@@ II Super-renormalizable Gravity
$S_{0}=\frac{2}{\kappa^{2}}\int\sqrt{-g}\,\mathrm{d}^{D}x\left(R+RF_{0}(\Box)R+R_{\mu\nu}F_{2}(\Box)R^{\mu\nu}+R_{\mu\nu\alpha\beta}F_{4}(\Box)R^{\mu\nu\alpha\beta}+V\right),$ (1)
where $R$ is the Riemann curvature scalar, $\kappa^{2}=32\pi G_{N}$, and a set of local terms $V$ cubic or higher in curvature. The latter involves operators with a carefully selected number of derivatives, ensuring that the favorable quantum properties of the theory are preserved. Furthermore, the crucial elements of our analysis are the functions of the covariant d’Alembertian operator $F_{i}(\Box)$ which called form factors. The form factors are assumed to be entire functions, allowing them to be expanded in a Taylor series $F_{i}(\Box)=\sum^{\infty}_{n=0}f_{in}\frac{\Box^{n}}{M^{2n}_{\ast}}$. where $M_{\ast}$ denotes the mass scale at which the higher-derivative terms in the action become relevant. Therefore, nonlocal gravity can be viewed as an infinite-order derivative extension of higher-derivative gravity. Such a model exhibits massive ghost particle excitations at the Planck scale $M_{P}$ and induces inflation through scalar particle excitations at the scale $m$ Stelle (1977); Koshelev et al. (2016); Ade and others (2016a, b). Consequently, the range for $M_{\ast}$ is naturally chosen as $m<M_{\ast}<M_{P}$. This phenomenological hierarchy suggests that the microscopical origins of inflation and nonlocal UV-completion of gravity are not the same.
Furthermore, the requirements of super-renormalizability and unitarity constrain the admissible form factors to the following categories
$\left\{\begin{split}&F_{0}(\Box)=-\frac{(D-2)\left(e^{H_{0}(\Box)}-1\right)+D\left(e^{H_{2}(\Box)}-1\right)}{4(D-1)\Box}+F_{4}(\Box),\\ &F_{2}(\Box)=\frac{e^{H_{2}(\Box)}-1}{\Box}-4F_{4}(\Box),\end{split}\right$. (2)
where $F_{4}(\Box)$ may in principle remain arbitrary, super-renormalizability mandates that $F_{4}(\Box)$ shares the same UV asymptotic behavior as the other two form factors $F_{i}(\Box)$ ($i=0,2$). To achieve this, the most straightforward approach is to retain only two of the three form factors, which leads to the condition $F_{4}(\Box)=0$. Additionally, the entire function $e^{H_{i}(\Box)}$ must satisfy three distinct sets of criteria Tomboulis (1997); Modesto (2015):
$\bullet$ The function $e^{H_{i}(\Box)}$ must be real and positive along the real axis and have no zeros within the entire complex plane for $z<\infty$ $\left(z\equiv-\frac{\Box}{M^{2}_{\ast}}\right)$. This requirement guarantees the absence of gauge-invariant poles, except for the transverse massless physical graviton pole.
$\bullet$ $e^{H_{i}(\Box)}$ exhibits the same asymptotic behavior along the real axis at $\pm\infty$.
$\bullet$ There exist value $0<\varphi<\frac{\pi}{2}$, and a positive integer $\gamma$, such that asymptotically
$|e^{H_{i}(\Box)}|\rightarrow|z|^{\gamma+N+1},\quad|z|\rightarrow\infty,\quad\text{with}\quad\gamma\geqslant\frac{D_{\text{even}}}{2}\quad\text{or}\quad\gamma\geqslant\frac{D_{\text{odd}}-1}{2},$ (3)
and regin $C$ defined by
$C\equiv\{z:-\varphi<\text{arg}z<+\varphi,\pi-\varphi<\text{arg}z<\pi+\varphi\}.$ (4)
The final condition is essential for achieving the maximum convergence of the theory in the UV regime. The required asymptotic behavior must be enforced not only along the real axis but also in the conical regions surrounding it. In Euclidean spacetime, condition (ii) is not strictly necessary if condition (iii) is satisfied. In Eq.(3), the capital $N$ is defined as a function of the spacetime dimension $D$, with $2N+4=D_{\text{odd}}+1$ for odd dimensions and $2N+4=D_{\text{even}}$ for even dimensions. Moreover, Ref.Tomboulis (1997) has already provided a specific example of a special form factor that satisfies the above conditions, which can be expressed as
$e^{H_{i}(\Box)}=e^{\frac{1}{2}\left[\Gamma(0,p_{i}(z)^{2})+\gamma_{E}+\log\left(p_{i}(z)^{2}\right)\right]},$ (5)
where $\gamma_{E}\approx 0.577216$ denotes the Euler-Mascheroni constant, $\Gamma(0,z)=\int_{z}^{\infty}\frac{e^{-t}}{t}\,\mathrm{d}t$ represents the incomplete Gamma function with its first argument set to zero. The polynomial $p_{i}(z)$, which has a degree $\gamma+N+1$ and satisfying $p_{i}(0)=0$, ensures that the low-energy limit of nonlocal theory is correct. In the UV regime, where $|z|\gg 1$, the function exhibits polynomial behavior of the form $|z|^{\gamma+N+1}$ within the conical region around the real axis, with an angular opening given by $\varphi=\frac{\pi}{4(\gamma+N+1)}$. To achieve super-renormalizability, it is necessary for the degrees of the polynomials defining $H_{0}(\Box)$ and $H_{2}(\Box)$ to be identical. Moreover, for simplicity, we will exclude the contributions from the interaction term $V$ in the subsequent analysis.
II.1 Propagator and Unitarity in Minkowski Spacetime
To obtain the explicit form of the propagator, we expand the metric tensor around Minkowski spacetime, denoted as $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$. In addition, we consider the following gauge-fixing term
$S_{\text{gf}}=\int\,\mathrm{d}^{D}x\frac{\eta^{\mu\nu}G_{\mu}\omega(\Box)G_{\nu}}{\kappa^{2}\xi},$ (6)
where $G_{\mu}$ is taken as $\partial^{\nu}h_{\mu\nu}$ to represent the usual harmonic gauge, $\xi$ denotes the gauge-fixing parameter, and $\omega(\Box)$ is a weight functional Accioly et al. (2002); Buchbinder et al. (1992); Stelle (1977). Subsequently, linearizing the action leads to the following expression
$S_{0}+S_{gf}\approx\frac{1}{2}\int\frac{\,\mathrm{d}^{D}q}{(2\pi)^{D}}\tilde{h}^{\mu\nu}(q)O_{\mu\nu\alpha\beta}\tilde{h}^{\alpha\beta}(-q),\\$ (7)
here we have already used the rescaled graviton field $\tilde{h}^{\mu\nu}\equiv\frac{h^{\mu\nu}}{\kappa}$. The operator $O$ is made out of two terms, one coming from the quadratization of Eq.(1) and the other from the gauge-fixing term. The d’Alembertian operator in $S_{0}$ and the gauge-fixing term are defined in flat spacetime. After inverting the operator $O$ Accioly et al. (2002) and applying the form factors from Eq.(LABEL:2), we derive the two-point function in momentum space,
@@ II.1 Propagator and Unitarity in Minkowski Spacetime
$\begin{split}O^{-1}_{\mu\nu\alpha\beta}&=\frac{1}{q^{2}e^{H_{2}(-q^{2})}}I_{\mu\nu\alpha\beta}+\left(\frac{2\xi}{q^{2}\omega(-q^{2})}-\frac{2}{q^{2}e^{H_{2}(-q^{2})}}\right)J_{\mu\nu\alpha\beta}\\ &-\left(\frac{4}{(D-1)(D-2)q^{2}e^{H_{0}(-q^{2})}}+\frac{4}{(D-1)q^{2}e^{H_{2}(-q^{2})}}\right)T_{\mu\nu\alpha\beta}\\ &+\left(\frac{2}{(D-1)q^{2}e^{H_{2}(-q^{2})}}-\frac{2}{(D-1)(D-2)q^{2}e^{H_{0}(-q^{2})}}\right)C_{\mu\nu\alpha\beta}\\ &+\left(\frac{D-2}{(D-1)q^{2}e^{H_{2}(-q^{2})}}+\frac{1}{(D-2)(D-1)q^{2}e^{H_{0}(-q^{2})}}-\frac{3\xi}{2q^{2}\omega(-q^{2})}\right)K_{\mu\nu\alpha\beta},\end{split}$ (8)
where the five independent operators are $I_{\alpha\beta}^{\mu\nu}=\frac{1}{2}\left(\delta_{\alpha}^{\mu}\delta_{\beta}^{\nu}+\delta_{\beta}^{\mu}\delta_{\alpha}^{\nu}\right)$$T_{\alpha\beta}^{\mu\nu}\equiv\frac{1}{4}\eta^{\mu\nu}\eta_{\alpha\beta}$, $C_{\alpha\beta}^{\mu\nu}\equiv\frac{1}{2}\left(\eta^{\mu\nu}\frac{q_{\alpha}q_{\beta}}{q^{2}}+\frac{q^{\mu}q^{\nu}}{q^{2}}\eta_{\alpha\beta}\right)$, $J_{\alpha\beta}^{\mu\nu}\equiv I_{\rho\kappa}^{\mu\nu}\frac{q_{\sigma}q^{\rho}}{q^{2}}I_{\alpha\beta}^{\sigma\kappa}$, and $K_{\alpha\beta}^{\mu\nu}\equiv\frac{q^{\mu}q^{\nu}}{q^{2}}\frac{q_{\alpha}q_{\beta}}{q^{2}}$. The propagator (8) represents the most general form consistent with unitarity. It only propagates the standard massless transverse spin-2 graviton, without introducing any additional d.o.f. This is due to the fact that exponents of entire functions are themselves special entire functions that do not have zeros, thus preventing the emergence of new poles and, consequently, any new physical d.o.f. Additionally, it is crucial to note that for the propagator to remain well-behaved, the correct forms of only two factors, corresponding to the $I_{\mu\nu\alpha\beta}$ and $T_{\mu\nu\alpha\beta}$ components, are required. This explains why one function out of three $F_{4}(\Box)$ can be put to zero from the point of view of unitarity. Furthermore, unitarity is explicitly preserved, as the optical theorem is trivially satisfied at the tree level, namely
$2\text{Im}\{T(q)^{\mu\nu}O^{-1}_{\mu\nu\alpha\beta}T(q)^{\alpha\beta}\}=2\pi\text{Res}\{T(q)^{\mu\nu}O^{-1}_{\mu\nu\alpha\beta}T(q)^{\alpha\beta}\}|_{q^{2}=0}>0,$ (9)
the $T^{\mu\nu}$ represents the energy-momentum tensor of matter field Koshelev et al. (2018). In particular, due to the conservation condition, it is sufficient to consider only the $I_{\mu\nu\alpha\beta}$ and $T_{\mu\nu\alpha\beta}$ components of the propagator.
II.2 Power counting in Minkowski Spacetime
We now revisit the power counting analysis of quantum divergences as presented in Refs.Kuzmin (1989); Tomboulis (1997, 2015); Eran (1998); Modesto and Rachwał (2015). In the UV regime, the propagator in momentum space exhibits the following scaling behavior
$O^{-1}\sim\frac{1}{q^{2\gamma+D}}.$ (10)
The vertices can be organized into different sets, some of which may involve the entire functions $e^{H_{\ell}(z)}$, while others may not. However, to establish a bound on the quantum divergences, it is sufficient to focus on the leading operators in the UV regime. These operators scale inversely with the propagator, thereby yielding upper bounds on the superficial degree of divergence for any graph $G$ Kuzmin (1989); Tomboulis (2015); Modesto (2012); Modesto and Rachwal (2014).
$\omega(G)=DL+(V-I)(2\gamma+D)$ (11)
in a spacetime of even or odd dimensionality $D$. Here, $V$ represents the number of vertices, $I$ denotes the number of internal lines, and $L$ stands for the number of loops. Furthermore, using the topological relation L1=IVL-1=I-V, we can simplify the expression as
$\omega(G)=D-2\gamma(L-1).$ (12)
Thus, if $\gamma>\frac{D}{2}$ (corresponding to Eq.(3)), only one-loop divergences remain in the theory, implying that higher-loop divergences are effectively suppressed. As a result, the theory is super-renormalizable Kuzmin (1989); Tomboulis (2015); Modesto (2012); Modesto and Rachwal (2014); Modesto and Rachwał (2015). In this case, only a finite number of operators, each with a mass dimension up to $D$, need to be included in the action to guarantee renormalizability in even-dimensional spacetime. This ensures that the theory remains well-defined at high energies, with a controlled and finite set of divergences, thus avoiding the issues associated with non-renormalizable theories.
III The Metric correction from Tree-Level Scattering Amplitudes
It is an intriguing concept that the Schwarzschild-Tangherlini metric can be derived from scattering amplitudes and Feynman diagrams Bjerrum-Bohr et al. (2018); Jakobsen (2020); Mougiakakos and Vanhove (2021). In this section, we examine the connection between super-renormalizable gravity and the vertex function of massive scalar particles interacting with gravitons. Specifically, we investigate whether it is possible to derive the gravitational metric from the scattering amplitudes. Besides, it is worth mentioning that our primary focus is on the classical limit of QG in flat spacetime, where, within this limit, the effects of ghost fields and divergent terms can be neglected. The classical limit of QFT is typically defined as the regime where $\hbar\to 0$. This is traditionally understood as the classical variational principle, $\delta S=0$. However, even in this limit, Feynman diagrams at any loop order can still make contributions. A detailed analysis of the classical limit of QFT can be found in Ref.Kosower et al. (2019), with specific attention to gravity in Refs.Bjerrum-Bohr et al. (2018); Cheung and Solon (2020); Bjerrum-Bohr et al. (2020). Furthermore, we consider two types of particles: massive scalar particles and gravitons. The massive scalars are treated as point particles, with their momenta remaining finite, while gravitons behave like waves in the classical theory, with their momenta approaching zero. These conclusions hold for both external and internal loop momenta. In this sense, the classical limit is analogous to a long-range limit.
The vertex rules come from the expansion of the matter term of the action which is